# Generative AI for Predicting 2D and 3D Wildfire Spread: Beyond Physics-Based Models and Traditional Deep Learning

### v1

@@ 5 State-of-the-Art Generative AI Applications in Wildfire Management
5.1 Fire Spread Prediction and Fire Behavior Modeling
@@
We identified seven existing studies that have explored or prototyped the use of various generative AI models to support fire spread prediction and fire behavior modeling. These studies leverage different architectures, including GANs, VAEs, and generative transformers, to simulate wildfire dynamics, generate synthetic fire scenarios, or enhance predictive performance. Collectively, they demonstrate the emerging potential of generative AI in capturing the complex spatiotemporal patterns of wildfire propagation and improving the accuracy, scalability, and adaptability of fire behavior forecasting systems.
5.2 Generative Adversarial Network (GAN)
@@
GANs and their variants have proven valuable in environmental and urban research by generating synthetic data to augment limited and imbalanced datasets (Xu et al., 2024c). This capability is especially critical in wildfire research, where collecting real-world fire data is often constrained by safety risks, high costs, and logistical challenges. By simulating underrepresented fire scenarios—such as passive and active crown fires—GANs help improve the diversity and robustness of training data, leading to more accurate classification of fire behavior and better support for operational decision-making. We identified two studies that effectively employ GANs for data augmentation to enhance wildfire spread prediction through integration with machine learning and simulation models.
@@
Khanmohammadi et al. (2023) presents a machine learning framework for predicting wildfire spread sustainability and crown fire occurrence in semiarid shrublands of southern Australia. The study uses a TGAN to generate synthetic fire records, addressing the challenge of small training datasets commonly encountered in this domain. The inclusion of synthetic data led to substantial performance improvements, increasing classification accuracy by up to 20% for spread sustainability and 4% for crown fire prediction. This enhanced dataset was used to train and evaluate several classifiers—including Support Vector Machines (SVM), Multilayered Neural Networks (MLP), and Multinomial Naive Bayes (MNB)—with the SVM model paired with TGAN-generated data achieving the highest accuracy. The results demonstrate the potential of combining generative AI with traditional supervised learning models to improve generalization and reliability in data-scarce wildfire prediction settings. Khanmohammadi et al. (2024) extends this approach to Canadian conifer forests, focusing on predicting wildfire propagation types, including surface, passive crown, and active crown fires. In this study, TGAN is again employed to address class imbalance by generating synthetic data for underrepresented fire types, significantly boosting prediction accuracy and F1-scores—especially for difficult-to-classify categories like passive crown fires. The study evaluates four additional machine learning models: two ensemble methods (Random Forest and XGBoost) and two AutoML approaches (TabPFN and AutoGluon), with TabPFN consistently achieving the best results when trained on GAN-augmented data. These findings reinforce the value of GAN-based data generation as a means of overcoming real-world data limitations and enhancing the predictive capabilities of wildfire behavior models.
@@ 5.2.1 Variational autoencoder (VAE)
Cheng et al. (2023) presents a generative AI framework aimed at improving spatial-temporal wildfire nowcasting by addressing the limitations of traditional physics-based simulation methods, which are often computationally intensive and time-consuming. The authors propose the use of a Vector-Quantized Variational Autoencoder (VQ-VAE) to generate high-fidelity, physically realistic sequences of burned area evolution in wildfire scenarios. The generative model is trained on synthetic data from a Cellular Automata (CA) simulator and used to produce 3D fire spread sequences with temporal and spatial consistency. This approach significantly accelerates data generation, achieving a speed-up of over four orders of magnitude compared to CA and MTT simulations—while capturing critical geophysical dependencies such as vegetation density and slope. The generated wildfire scenarios serve as training data for a machine learning surrogate model combining Proper Orthogonal Decomposition (POD) and a Long Short-Term Memory (LSTM) network to emulate fire dynamics. The use of VQ-VAE addresses key challenges in the domain, including the high computational cost of simulation, the need for large volumes of training data, and the ability to simulate new, realistic wildfire events that reflect observed behavior. The inclusion of VQ-VAE–generated data improved the surrogate model’s prediction accuracy and structural similarity scores on both simulated and real wildfire events, demonstrating the effectiveness of generative AI in enhancing the resolution, efficiency, and generalizability of wildfire forecasting systems.
@@ 5.2.2 Transformer
Li and Rad (2024) presents a transformer-based hybrid model—Attention Swin U-Net with Focal Modulation (ASUFM)—designed for next-day wildfire spread forecasting across North America. ASUFM integrates spatial attention and focal modulation layers into a Swin Transformer U-Net backbone, producing predictive fire masks from multivariate remote sensing data. Though not a conventional generative model, ASUFM simulates realistic fire spread scenarios and addresses challenges such as spatial resolution, class imbalance, and temporal consistency. Trained on an expanded NDWS dataset (2012–2023), the model achieved state-of-the-art results in Dice score and PR-AUC, outperforming U-Net and other transformer variants. Additional techniques, including weighted loss functions, skip connections, and cosine learning rate scheduling, further enhanced the model’s generalization and accuracy. Deepa et al. (2024) proposes a contrastive learning framework for early forest fire prediction that combines a Contrastive Vision Transformer (CViT) with a Pool Former module. CViT functions as a powerful feature extractor via self-supervised contrastive learning and multi-head attention, while the Pool Former improves prediction efficiency by modeling spatial dependencies without heavy matrix operations. Though it does not employ traditional generative AI, the framework enhances feature robustness under variable environmental conditions through preprocessing and data augmentation strategies. The model achieved competitive performance, with 92.8% accuracy and an F1-score of 90.2%, surpassing multiple baseline models including CNN, RNN, Faster R-CNN, and ResNet. Annane et al. (2024) introduces a real-time wildfire prediction system that integrates a CNN–Transformer hybrid model with a blockchain-based infrastructure for enhanced data security and traceability. The CNN extracts spatial features from drone imagery, while the transformer captures temporal patterns such as smoke visibility and fire direction. Although it does not use generative AI in the traditional sense, the system generates fire probability maps that simulate the spread and severity of wildfires, improving alert accuracy and resolution. It also incorporates a rule-based assistant for decision support and achieves 93.18% prediction accuracy, outperforming CNN and ResNet-42 models, albeit with a slightly higher training time. In contrast, Li et al. (2024b) presents a true generative AI approach through the development of Sim2Real-Fire, a large-scale, multi-modal dataset paired with the S2R-FireTr transformer model. S2R-FireTr forecasts and backtracks binary fire masks by learning from 1 million simulated wildfire sequences and generalizing to 1,000 real-world cases. By leveraging cross-attention across five aligned modalities—topography, vegetation, fuel types, weather, and satellite imagery—it generates physically plausible fire spread scenarios even under temporally incomplete conditions. Outperforming physical simulators and deep learning baselines in both AUPRC (72.9%) and F1-score (69.6%), S2R-FireTr significantly reduces simulation time while maintaining high spatial-temporal fidelity. This work exemplifies the transformative potential of generative AI in enabling rapid, scalable, and low-cost wildfire scenario generation for both predictive and forensic applications.
5.3 Wildfire Detection and Monitoring using Vision Transformers
@@
Generative AI models, particularly vision transformers, are increasingly utilized in wildfire monitoring to generate high-resolution segmentation masks and predictive spatial representations of fire spread. These models excel in tasks requiring fine-grained spatial understanding and temporal consistency, such as detecting smoke plumes and segmenting active fire zones, due to the visual complexity of input data (e.g., RGB UAV imagery, thermal maps) and the operational need for accurate, real-time predictions to support early firefighting efforts. Our review identified three notable studies that incorporate transformer-based architectures to facilitate wildfire detection and monitoring.
@@
Falcão et al. (2023) focuses on early detection of wildfire smoke plumes using a deep learning ensemble framework that integrates EfficientNetV2 (CNN), DeiT, and Swin TransformerV2 (both vision transformers). While generative models are not explicitly employed, the ensemble architecture—coupled with a neural network-based meta-classifier—demonstrates strong performance under challenging conditions like haze, fog, and low-contrast smoke. The study leverages transfer learning and data augmentation techniques to improve generalization, achieving an average accuracy of 96.46% and an AUPRC of 95.14%, highlighting the value of architectural diversity in image-based wildfire surveillance. Ghali et al. (2022) targets fire classification and segmentation using UAV imagery. Although it does not incorporate conventional generative AI, the framework produces fine-grained segmentation masks, effectively achieving a generative outcome. The authors develop an ensemble classifier with EfficientNet-B5 and DenseNet-201 for fire detection and employ three segmentation models—EfficientSeg, TransUNet, and TransFire—two of which are transformer-based. These models address complex challenges such as detecting small fires and delineating fire boundaries in noisy, cluttered backgrounds. TransUNet-R50-ViT outperforms CNN-based baselines on the FLAME UAV dataset, achieving an F1-score of 99.9%, confirming the efficacy of vision transformers in high-resolution fire segmentation. Ghali et al. (2021) explores the use of vision transformers—TransUNet and Medical Transformer (MedT)—to segment wildfire regions from RGB imagery for early detection and boundary delineation. These transformers perform generative tasks by producing binary segmentation masks that can be used to simulate wildfire spread scenarios. The models address core challenges such as fine boundary detection, long-range dependency modeling, and feature misclassification. Trained on the CorsicanFire dataset, TransUNet and MedT achieve F1-scores of 97.7% and 96.0% respectively, outperforming conventional architectures like U-Net, U2-Net, and EfficientSeg. Despite slightly longer inference times (1.2s for TransUNet and 2.72s for MedT), the trade-offs are justified by improvements in spatial granularity, generalization, and robustness under diverse environmental conditions.
5.4 Wildfire Risk Mapping using Transformers

### v2

@@ 5 Generative AI Applications in Wildfire Management
5.1 Fire Spread Prediction
@@
We identified seven existing studies that have explored or prototyped the use of various generative AI models to support fire spread prediction and fire behavior modeling. These studies leverage three different architectures, namely GANs, VAEs, and generative transformers, to simulate wildfire dynamics, generate synthetic fire scenarios, or enhance predictive performance. Collectively, they demonstrate the emerging potential of generative AI in capturing the complex spatiotemporal patterns of wildfire propagation and improving the accuracy, scalability, and adaptability of fire behavior forecasting systems.
5.1.1 Generative Adversarial Network (GAN)
@@
GANs and their variants have proven valuable in environmental and urban research by generating synthetic data to augment limited and imbalanced datasets Xu et al. (2024c). This capability is especially critical in wildfire research, where collecting real-world fire data is often constrained by safety risks, high costs, and logistical challenges. By simulating underrepresented fire scenarios—such as passive and active crown fires—GANs help improve the diversity and robustness of training data, leading to more accurate classification of fire behavior and better support for operational decision-making. We identified two studies that effectively employ GANs for data augmentation to enhance wildfire spread prediction through integration with machine learning and simulation models.
@@
Khanmohammadi et al. (2023) presents a machine learning framework for predicting wildfire spread sustainability and crown fire occurrence in semiarid shrublands of southern Australia. The study uses a TGAN to generate synthetic fire records, addressing the challenge of small training datasets commonly encountered in this domain. The inclusion of synthetic data led to substantial performance improvements, increasing classification accuracy by up to 20% for spread sustainability and 4% for crown fire prediction. This enhanced dataset was used to train and evaluate several classifiers—including Support Vector Machines (SVM), Multilayered Neural Networks (MLP), and Multinomial Naive Bayes (MNB)—with the SVM model paired with TGAN-generated data achieving the highest accuracy. The results demonstrate the potential of combining generative AI with traditional supervised learning models to improve generalization and reliability in data-scarce wildfire prediction settings. Trained on a dataset comprising 61 experimental fires in Australian semiarid shrublands—augmented by 27 TGAN-generated samples for spread sustainability and 13 for crown fire cases—the TGAN-enhanced SVM achieved 90% accuracy for predicting sustained fire propagation and 80% accuracy for active crown fire occurrence on an independent evaluation set of 29 fires. Compared to logistic regression models developed on the same data, this represented improvements of 15 percentage points for spread sustainability and 4 points for crown fire prediction, along with markedly higher sensitivities (1.0 vs. 0.77 and 0.55, respectively). Although the study did not report training or inference times, it clearly demonstrated that integrating TGAN-based data augmentation substantially boosts predictive performance and robustness relative to traditional statistical approaches, especially under data-limited conditions.
Khanmohammadi et al. (2024) extends this approach to Canadian conifer forests, focusing on predicting wildfire propagation types, including surface, passive crown, and active crown fires. In this study, TGAN is again employed to address class imbalance by generating synthetic data for underrepresented fire types, significantly boosting prediction accuracy and F1-scores—especially for difficult-to-classify categories like passive crown fires. The study evaluates four additional machine learning models: two ensemble methods (Random Forest and XGBoost) and two AutoML approaches (TabPFN and AutoGluon), with TabPFN consistently achieving the best results when trained on GAN-augmented data. These findings reinforce the value of GAN-based data generation as a means of overcoming real-world data limitations and enhancing the predictive capabilities of wildfire behavior models. Trained on a dataset of 113 experimental fire cases collected over 60 years in Canadian conifer forests, including 52 surface fires, 23 passive crown, and 38 active crown fires, the study demonstrated that using TGAN to generate synthetic records dramatically improved prediction performance. With balanced datasets created by adding up to 26 synthetic passive crown and 13 active crown fire samples, the best generative AI-related model, TabPFN, achieved 91% accuracy and a 93% F1-score on independent evaluation data in the binary classification task, outperforming traditional logistic regression baselines. In the more challenging multi-class setting (surface vs. passive crown vs. active crown), the addition of GAN data elevated TabPFN’s accuracy from 61% to 74% and its F1-score from 50% to 65%, particularly improving predictions of passive crown fires. Although exact training or inference times were not provided, the study clearly highlighted how TGAN-based data augmentation substantially enhances the spatial-temporal prediction fidelity of wildfire models under data-limited scenarios.
In Summary, previous studies have shown that integrating GAN-based data augmentation can substantially outperform traditional simulation and machine learning approaches in fire spread prediction and behavior modeling Khanmohammadi et al. (2023, 2024). By leveraging TGANs to generate synthetic wildfire records, these works effectively addressed class imbalance and limited data, achieving up to 90–91% accuracy and F1-scores exceeding 93%, notably outperforming logistic regression baselines. In more complex multi-class settings involving surface, passive crown, and active crown fires, GAN-augmented models improved accuracy by over 13 percentage points and F1-scores by 15 points, significantly enhancing the detection of underrepresented fire types. These findings highlight the strength of GAN-driven data generation in boosting the fidelity, robustness, and generalizability of wildfire forecasting models beyond what conventional statistical or standard deep learning methods can achieve under data-scarce conditions.
@@ 5.1.2 Variational autoencoder (VAE)
Cheng et al. (2023) introduces a generative AI framework to overcome the computational demands of traditional physics-based wildfire simulations. Using a Vector-Quantized Variational Autoencoder (VQ-VAE) trained on Cellular Automata (CA) data, the model generates high-fidelity 3D sequences of wildfire spread that capture essential geophysical influences like vegetation and slope. This approach accelerates data generation by over four orders of magnitude compared to CA and MTT simulators. The synthetic wildfire scenarios are then used to train a POD-LSTM surrogate model, which, with the inclusion of VQ-VAE data, achieves notably higher prediction accuracy and structural similarity on both simulated and real events. This demonstrates the VQ-VAE’s effectiveness in reducing simulation costs, enriching training datasets, and improving the realism and generalizability of wildfire forecasts. Specifically, the VQ-VAE was trained on 40 CA-generated fire spread sequences from the Chimney Fire region in California, each spanning 8 days (16 temporal snapshots) at 128×128 spatial resolution, and was then used to generate 500 additional synthetic fire scenarios. This resulted in surrogate models trained on VQ-VAE–augmented data achieving significantly lower relative RMSE and higher structural similarity index (SSIM) than models trained solely on the limited CA data, for both unseen synthetic fires and the real Chimney Fire event observed via MODIS and VIIRS satellites. In terms of computational efficiency, the VQ-VAE generated 8-day wildfire spread sequences in just 0.3 seconds, representing a speed-up of 4–5 orders of magnitude compared to traditional CA and MTT simulators, effectively overcoming the major runtime bottlenecks of physics-based wildfire modeling while preserving crucial spatial-temporal dynamics.
@@ 5.1.3 Transformer
Li and Rad (2024) presents a transformer-based hybrid model—Attention Swin U-Net with Focal Modulation (ASUFM)—designed for next-day wildfire spread forecasting across North America. ASUFM integrates spatial attention and focal modulation layers into a Swin Transformer U-Net backbone, producing predictive fire masks from multivariate remote sensing data. Though not a conventional generative model, ASUFM simulates realistic fire spread scenarios and addresses challenges such as spatial resolution, class imbalance, and temporal consistency. Trained on an expanded NDWS dataset (2012–2023), the model achieved state-of-the-art results in Dice score and PR-AUC, outperforming U-Net and other transformer variants. Additional techniques, including weighted loss functions, skip connections, and cosine learning rate scheduling, further enhanced the model’s generalization and accuracy. On the extended NDWS dataset covering over 31,000 samples from 2012–2023 across all of North America at 1 km spatial resolution, ASUFM achieved a Dice score of 0.4066, precision of 0.4345, recall of 0.4096, and a PR-AUC of 0.3974, substantially outperforming traditional encoder-decoder baselines such as U-Net (Dice 0.3493, PR-AUC 0.2945) and even pure Swin U-Net variants. This highlights the transformer’s superior ability to model long-range dependencies and capture complex spatial fire propagation patterns. Although the study did not report explicit training or inference times, it noted using large-scale GPU infrastructure (NVIDIA A100-80G or RTX A6000) with advanced techniques like focal modulation and spatial attention to balance precision-recall under heavy class imbalance. The expanded spatial and temporal coverage of the dataset ensures that ASUFM’s learned representations generalize across diverse geographic terrains and multi-year fire regimes, demonstrating its promise for continent-scale, next-day wildfire spread prediction.
Deepa et al. (2024) proposes a contrastive learning framework for early forest fire prediction that combines a Contrastive Vision Transformer (CViT) with a Pool Former module. CViT functions as a powerful feature extractor via self-supervised contrastive learning and multi-head attention, while the Pool Former improves prediction efficiency by modeling spatial dependencies without heavy matrix operations. Though it does not employ traditional generative AI, the framework enhances feature robustness under variable environmental conditions through preprocessing and data augmentation strategies. The study utilized the publicly available Forest Fire Big Data dataset from Kaggle, comprising 1,832 images of forest scenes under diverse environmental and weather conditions, covering both fire and no-fire instances. In terms of quantitative performance, the CViT-Pool Former method achieved 92.8% accuracy, 89.4% precision, 91.4% recall, and a 90.2% F1-score, clearly outperforming Faster R-CNN (accuracy 89.0%, F1 88.9%) and ResNet (accuracy 87.4%, F1 84.5%). While the study did not report explicit training or inference times, it demonstrated that transformer-based contrastive learning significantly boosts predictive robustness and spatial discrimination compared to traditional deep CNN and RNN baselines.
Annane et al. (2024) introduces a real-time wildfire prediction system that integrates a CNN–Transformer hybrid model with a blockchain-based infrastructure for enhanced data security and traceability. The CNN extracts spatial features from drone imagery, while the transformer captures temporal patterns such as smoke visibility and fire direction. Although it does not use generative AI in the traditional sense, the system generates fire probability maps that simulate the spread and severity of wildfires, improving alert accuracy and resolution. It also incorporates a rule-based assistant for decision support and achieves 93.18% prediction accuracy, outperforming CNN and ResNet-42 models, albeit with a slightly higher training time. The system was trained and tested on forest fire images from Algeria’s Béjaïa region, comprising 540 training images and 100 test images, each at a spatial resolution of 244×244 pixels. In quantitative terms, the CNN–Transformer model achieved an accuracy of 93.18%, precision of 91%, recall of 97%, and an F1-score of 94%, surpassing both standalone CNN and CNN ResNet-18 baselines. While it required a training time of approximately 30 minutes—slightly longer than the 20 minutes for CNN ResNet-18—this additional computational cost was offset by substantially higher predictive performance, particularly in modeling complex temporal features like fire evolution and direction. These results highlight the efficacy of integrating transformer-based temporal learning in enhancing the spatial-temporal resolution of wildfire risk maps.
In contrast, Li et al. (2024b) presents a true generative AI approach through the development of Sim2Real-Fire, a large-scale, multi-modal dataset paired with the S2R-FireTr transformer model. S2R-FireTr forecasts and backtracks binary fire masks by learning from 1 million simulated wildfire sequences and generalizing to 1,000 real-world cases. By leveraging cross-attention across five aligned modalities—topography, vegetation, fuel types, weather, and satellite imagery—it generates physically plausible fire spread scenarios even under temporally incomplete conditions. Trained on 1 million multi-modal simulated wildfire scenarios with global coverage at 30 m spatial resolution, and evaluated on 1,000 annotated real-world wildfire cases, S2R-FireTr achieved an AUPRC of 72.9%, F1-score of 69.6%, and IoU of 56.4% on real-world data. These results substantially outperformed traditional physical simulators—FARSITE (AUPRC 55.9%), WFDS (61.2%), and WRF-SFIRE (63.0%)—as well as leading deep learning baselines like Rainformer and Earthformer. While exact training or inference times were not detailed, the approach demonstrated orders-of-magnitude efficiency gains by bypassing computationally intensive physics-based simulations, reinforcing its value for large-scale, real-time wildfire forecast and backtracking tasks.
Together, these studies pave the way for broader adoption of advanced transformer-based and generative AI methods, which have likewise shown remarkable improvements over traditional simulation tools and conventional deep learning in wildfire forecasting. For example, Li and Rad (2024) reported that ASUFM, trained on 31,000 samples across North America, achieved superior Dice and PR-AUC metrics compared to U-Net and Swin U-Net, effectively modeling complex spatial fire dynamics at continental scale. Similarly, Deepa et al. (2024) found that a contrastive vision transformer pipeline outperformed CNN and RNN baselines, reaching over 92% accuracy and a 90% F1-score on diverse forest imagery. Annane et al. (2024) further demonstrated that integrating CNNs with transformers achieved 93% accuracy and a 94% F1-score on Algerian wildfire data, capturing temporal evolution beyond what CNNs alone could model. Finally, Li et al. (2024b) showed that a transformer trained on 1 million globally simulated wildfire scenarios attained AUPRC gains of 9–17 points over physics-based models like FARSITE and WFDS, enabling near real-time large-scale forecasting. Collectively, these results underscore the transformative advantages of GANs, VAEs, and transformers in enhancing the precision, scalability, and efficiency of wildfire spread prediction and behavior modeling.
5.2 Wildfire Detection and Monitoring
@@
Vision transformers are increasingly utilized in wildfire monitoring to generate high-resolution segmentation masks and predictive spatial representations of fire spread. These models excel in tasks requiring fine-grained spatial understanding and temporal consistency, such as detecting smoke plumes and segmenting active fire zones, due to the visual complexity of input data (e.g., RGB UAV imagery, thermal maps) and the operational need for accurate, real-time predictions to support early firefighting efforts. Our review identified three notable studies that incorporate transformer-based architectures to facilitate wildfire detection and monitoring.
Falcão et al. (2023) focuses on early detection of wildfire smoke plumes using a deep learning ensemble framework that integrates EfficientNetV2 (CNN), DeiT, and Swin TransformerV2 (both vision transformers). While generative models are not explicitly employed, the ensemble architecture—coupled with a neural network-based meta-classifier—demonstrates strong performance under challenging conditions like haze, fog, and low-contrast smoke. The study leverages transfer learning and data augmentation techniques to improve generalization, achieving an average accuracy of 96.46% and an AUPRC of 95.14%, improving upon the best individual base model by approximately 2.1% in accuracy and 1.2% in AUPRC. Precision remained comparably high to the strongest base learner (EfficientNetV2), while recall and F1-scores demonstrated consistent gains. Although the study did not benchmark against traditional machine learning or classical computer vision methods, it provided clear computational insights—training was performed via fine-tuning with transfer learning on a Google Colab platform using NVIDIA T4 GPUs, and inference time was measured at approximately 58 ms per batch of 16 images, compared to around 17 ms for individual models. This moderate increase in latency was justified by substantially enhanced robustness, allowing real-time deployment across hundreds of camera feeds operating at practical frame rates for wildfire monitoring.
Ghali et al. (2022) targets fire classification and segmentation using UAV imagery. Although it does not incorporate conventional generative AI, the framework produces fine-grained segmentation masks, effectively achieving a generative outcome. The authors develop an ensemble classifier with EfficientNet-B5 and DenseNet-201 for fire detection and employ three segmentation models—EfficientSeg, TransUNet, and TransFire—two of which are transformer-based. These models address complex challenges such as detecting small fires and delineating fire boundaries in noisy, cluttered backgrounds. The proposed transformer-based segmentation models demonstrated remarkable quantitative gains. TransUNet-R50-ViT achieved an accuracy and F1-score of 99.9%, outperforming the CNN-based U-Net by nearly 1 percentage point, and effectively capturing fine fire boundaries even in cluttered backgrounds. TransFire also excelled with an F1-score of 99.82%, confirming the advantage of vision transformers in extracting detailed fire regions. In terms of computational performance, TransUNet completed inference in 0.51 s per image, slightly higher than U-Net’s 0.29 s, while TransFire processed images at 1.0 s, reflecting a modest increase in latency for substantially improved segmentation fidelity. Although the study did not benchmark these models against traditional machine learning approaches, it clearly demonstrated the superior precision and robustness of transformer-based architectures in high-resolution wildfire segmentation using UAV imagery.
@@
Ghali et al. (2021) explores the use of vision transformers—TransUNet and Medical Transformer (MedT)—to segment wildfire regions from RGB imagery for early detection and boundary delineation. These transformers perform generative tasks by producing binary segmentation masks that can be used to simulate wildfire spread scenarios. The models address core challenges such as fine boundary detection, long-range dependency modeling, and feature misclassification. Trained on the CorsicanFire dataset, TransUNet and MedT achieve F1-scores of 97.7% and 96.0% respectively, outperforming conventional architectures like U-Net, U2-Net, and EfficientSeg. Despite slightly longer inference times (1.2s for TransUNet and 2.72s for MedT), the trade-offs are justified by improvements in spatial granularity, generalization, and robustness under diverse environmental conditions. In terms of quantitative performance, TransUNet achieved an F1-score of 97.7% with its hybrid ResNet50-ViT backbone, while MedT reached 96.0%, each substantially exceeding the benchmarks set by U-Net (92.0%), U2-Net (82.9%), and EfficientSeg (94.3%). Precision and recall were jointly reflected in these high F1-scores, underscoring the transformers’ capability to minimize both false positives and false negatives in fire boundary segmentation. Although the study did not report training time in detail, inference times were noted at 1.2 seconds for TransUNet and 2.72 seconds for MedT, representing a moderate computational overhead compared to simpler models, yet yielding superior segmentation fidelity and finer detection of small or diffuse fire regions.
In terms of performance, transformer-based models have consistently outperformed traditional simulation tools and conventional deep learning approaches in both fire spread detection and boundary segmentation. For instance, Falcão et al. (2023) demonstrated that stacking vision transformers with CNNs boosted average accuracy to 96.46% and AUPRC to 95.14%, surpassing the best individual models by notable margins while maintaining practical inference times suitable for real-time monitoring. Similarly, Ghali et al. (2022) and Ghali et al. (2021) showed that transformer-based segmentation networks like TransUNet and TransFire achieved exceptional F1-scores up to 99.9%, outperforming CNN-based baselines such as U-Net by 5–15 percentage points, with only modest increases in inference time. These results highlight the superior ability of transformers to capture fine spatial details and long-range dependencies, enhancing the precision, robustness, and spatial fidelity of wildfire detection and boundary modeling beyond what traditional or purely CNN-based methods can achieve.
5.3 Wildfire Risk Mapping